# ESG-Net: Event-Aware Semantic Guided Network for Dense Audio-Visual Event Localization

### v1

@@ III-C Early Semantics Interaction
(3) Cross-Modal Pyramid. Following [1], we perform futher cross-modal fusion to enable the model to better understand audio-visual events at different temporal resolutions. Specifically, we temporally downsample the features $\{F_{AV},F_{VA}\}$ with the stride $2^{l_{c}-1}$, where $l_{c}$ is the index of the current pyramid block. There is a cross-attention after a downsampling of audio and visual features. The outputs of the cross-modal pyramid module are $\{\hat{F}_{AV},\hat{F}_{VA}\}\in\mathbb{R}^{T_{l}\times D}$, where $T_{l}$ is the total length of the multi-level pyramid features.
Multi-Stage Semantic Guidance. In this section, we aim to enhance audio-visual event localization by enabling the model to understand event-related content in intermediate layers. However, even with the Multi-Modal Early Fusion, the model still struggles to understand the principles of audio-visual event localization clearly. Due to the significant differences between audio and video features, intermediate layers without cross-modal semantic bridging may cause modality semantic gap. Inspired by [28, 56], we apply an explicit event-related semantic guidance that directly facilitate audio-visual events understanding through multi-stage constraints. Specifically, we employ MLPs on the outputs of Multi-Modal Early Fusion ($\{\hat{F}_{A},\hat{F}_{V}\}$, $\{F_{AV},F_{VA}\}$, $\{\hat{F}_{AV},\hat{F}_{VA}\}$) for decoding. Each MLP consists of two linear layers with a PReLU activation. Then, we use focal loss functions [57] respectively on the outputs after MLPs for semantic consistency constraints. Note that here we only detect the event categories for each snippet, without regressing the temporal boundaries. The total multi-stage classification loss $\mathcal{L}_{mcls}$ is as:
$\mathcal{L}_{mcls}=\alpha_{1}\mathcal{L}_{AV1}+\alpha_{2}\mathcal{L}_{AV2}+\alpha_{3}\mathcal{L}_{AV3}$ (8)
$\mathcal{L}_{AVi}=\mathcal{L}_{Ai}+\mathcal{L}_{Vi}$ (9)
where $\alpha_{i}$ is the weight of losses, $i\in\{1,2,3\}$. $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ increase sequentially to ensure the effectiveness of multi-stage optimization from single-modal to multi-modal information.
@@ III-D Mixture of Dependency Experts
Multi-Modal Temporal Aggregation. Before localizing audio-visual events, we need to aggregate the results of ESI $\{\hat{F}_{AV},\hat{F}_{VA}\}$ and apply temporal modeling to obtain a comprehensive multi-modal representation. By doing this, the model can capture cross-modal long-term relations and filter out irrelevant noise. Specifically, $\hat{F}_{AV}$ and $\hat{F}_{VA}$ are first concatenated together as $Z$. At this point, $Z$ has already contained comprehensive multi-modal information, and has undergone multi-stage semantic guidance. However, $\hat{F}_{AV}$ and $\hat{F}_{VA}$ contain some redundant information, so we first input $Z$ into a projection layer to get $Z^{\prime}\in\mathbb{R}^{T_{l}\times C}$ for preliminary fusion, where $C$ is dimension of $Z^{\prime}$, and it is also the number of event categories. Then, to further refine multi-modal information, we employ a temporal modeling for $Z^{\prime}$ as in Eq. (1) and Eq. (2), composed of $N_{1}+N_{2}$ multi-headed self-attention. $\tilde{Z}\in\mathbb{R}^{T_{l}\times C}$is the output after $N_{1}$ temporal attention blocks, and $Z_{t}$ is the final output after $N_{1}+N_{2}$ temporal attention blocks.
Multi-Event Dependencies Extraction. In audio-visual scenarios, there are dependencies between concurrent events (e.g., thunder is often accompanied by rain and wind). However, these correlations are complex and variable across different scenarios, and the events may have different degrees of overlap. We hope that the model can reveal event dependencies automatically according to different inputs. We employ a series of MoE [8, 9, 10] layers as in the lower right part of Figure 2. Each MoE layer consists of $n$ experts, with only one expert selected at a time. By combining different experts across $m$ layers, complex multi-event dependencies can be revealed adaptively (each MoE layer has $n$ experts, there would be $n^{m}$ possible combinations). Specifically, to ensure that only one expert is selected in each layer at a time, we designed the MoE gate with a Gumbel-Softmax [58]. The output of $i$-th layer is:
@@
where $E_{i,j}$ is the $j$-th expert in the $i$-th layer, $Z_{e}^{i-1}\in\mathbb{R}^{T_{l}\times C}$ is the output of $(i-1)$-th MoE layer. Finally, the output of the last MoE layer $Z_{e}$ is added with $Z_{t}$ to obtain $\hat{Z}$ for decoding.
@@ III-E Training and Inference
Training. We employ an end-to-end training approach, where three losses are used to constrain the following aspects: a focal loss [57] $\mathcal{L}_{cls}$ for event classification, a generalized IoU loss [59] $\mathcal{L}_{reg}$ for temporal boundary estimation, and $\mathcal{L}_{mcls}$ for multi-stage semantic guidance classification as in Eq. (8) and Eq. (9). The total loss function is expressed as:
@@
$\mathcal{L}_{cls}$ and $\mathcal{L}_{reg}$ are used to constrain the final audio-visual event detection and regression, while $\mathcal{L}_{mcls}$ is used for cross-modal semantic bridging in intermediate layers. Note that features without sufficient cross-modal modeling cannot regress accurate temporal boundaries of the event, so $\mathcal{L}_{mcls}$ does not include the generalized IoU loss for boundry regression.
Inference. When given all snippets of a video during inference, the model will output candidates that include event categories and temporal boundaries, each accompanied by a confidence score. Then, all candidates are processed by multi-class Soft-NMS [60] to filter out temporal boundaries with high overlap within the same event category. For convenience, the whole process of ESG-Net is summarized in Algorithm 1.
Algorithm 1 Proposed ESG-Net.
  Input: Paired audio and visual snippets $A=\{A_{t}\}_{t=1}^{T}$, $V=\{V_{t}\}_{t=1}^{T}$ with GT $Y=\{c_{n},t^{s}_{n},t^{e}_{n}\}_{n=1}^{N}$, hyper-parameters $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ and $\{m,n\}$, maximum iterations MaxEpoch.
  Output: Prediction result $\hat{P}=\{\hat{c}_{n},\hat{t}^{s}_{n},\hat{t}^{e}_{n}\}_{n=1}^{N}$ and trainable parameters of ESG-Net.
  Initialization: Initialize network parameters of ESI module, MoDE module and Decoder. Load pre-trained parameters of audio and visual encoder $\mathscr{E}_{A}$ and $\mathscr{E}_{V}$.
  Training
     for epoch = 0 to MaxEpoch do
        1. Extract audio features ${F_{A}}$ and visual features ${F_{V}}$       by frozen encoders $\mathscr{E}_{A}$ and $\mathscr{E}_{V}$.
        2. Perform Multi-Modal Early Fusion to obtain       $\{\hat{F}_{A},\hat{F}_{V}\}$, $\{F_{AV},F_{VA}\}$ and $\{\hat{F}_{AV},\hat{F}_{VA}\}$.
        3. Compute the multi-stage semantic guidance loss       $\mathcal{L}_{mcls}$ by Eq. (8), (9).
        4. Perform MoDE to obtain $\hat{Z}$.
        5. Compute predicted results $\hat{P}=\{\hat{c}_{n},\hat{t}^{s}_{n},\hat{t}^{e}_{n}\}_{n=1}^{N}$.
        6. Compute the total loss $\mathcal{L}$ by Eq. (13)
        7. Update trainable network parameters to minimize $\mathcal{L}$.
     end for
  Inference
     Input test samples and obtain $\hat{P}=\{\hat{c}_{n},\hat{t}^{s}_{n},\hat{t}^{e}_{n}\}_{n=1}^{N}$.
@@ IV-A Experimental Settings
Datasets. The UnAV-100 dataset proposed by Geng et al. [1] consists of 10,790 videos with a total duration exceeding 126 hours, and includes 30,059 audio-visual events across 100 event categories. In addition to the raw videos, UnAV-100 also provides audio and video features extracted by frozen encoders. It includes two versions: (1) The optical flow features are extracted by RAFT [61], the RGB features are extracted by two-stream I3D [11], while audio features are extracted by VGGish [12]; (2) Audio and video features are extracted by ONE-PEACE [13], where the visual encoder is fine-tuned on Kinetics-400 [62]. As the standard dataset distribution, the training, validation and testing set are split at a ratio of 3:1:1.
Baseline Models. We first compare several state-of-the-art TAL [21] models, including VSGN [63], TadTR [22], ActionFormer [23] and TridDet [64]. These methods were used for trimmed videos while each video contains only a single event. We also compare the baseline method for DAVE, UnAV [1], which is the first audio-visual event localization method applied to untrimmed videos with dense events. Additionally, we compare the existing state-of-the-art DAVE models, UniAV [5], LoCo [6] and CCNet [33]. As shown in Table I and Table II, UniAV(AT) denotes the model trained jointly on TAL [21], SED [24], and AVEL [14, 15] tasks, while UniAV(STF) denotes the model that is further fine-tuned for AVEL based on UniAV(AT).
@@
Implementation Details. As with the previous methods [1, 5], the inputs settings of two versions (I3D-VGG [11, 12] and ONE-PEACE [13]) are as follows: (1) the sampling rates of video frames are 25 fps and 16 fps respectively, the sizes of sliding windows are 24 frames and 16 frames with the strides of 8 and 4, while each audio is sampled in alignment with the video; (2) the maximum sequence lengths are $\{T_{1}=224$, $T_{2}=256\}$ by cropping or padding, the dimensions of (audio and video) features are $\{D_{A1}=128,D_{V1}=2048\}$ and $\{D_{A2}=1536,D_{V2}=1536\}$, respectively. The embedding dimension in our model is 512, $L_{c}=6$, $N_{1}=2$, and $N_{2}=1$. The layers of MoDE is $m=4$, while the number of experts in each layer is $n=2$. The weights in $\mathcal{L}_{mcls}$ are set to $\alpha_{1}=0.3$, $\alpha_{2}=0.6$ and $\alpha_{3}=0.9$. We use Adam optimizer [65] to train our model for 40 epochs with a linear warmup of 5 epochs. Specifically, the batch size is 9, the initial learning rate is 1e-3 and the weight decay is 1e-4. Our expertments are implemented in an end-to-end manner on one RTX 3090 GPU with a 24GB memory.
@@ IV-B Comparison with Existing Work
As shown in Table I and Table II, we conduct comprehensive performance comparisons between the proposed method and existing methods. These results show that ESG-Net outperforms other methods in overall performance when using I3D+VGGish [11, 12] or ONE-PEACE [13] for features extracting. Compared to the existing state-of-the-art TAL models (VSGN [63], TadTR [22], ActionFormer [23] and TridDet [64]), ESG-Net demonstrates significant performance improvements at any threshold, indicating that our method is suitable for untrimmed videos with dense audio-visual events. When compared to the baseline method (UnAV [1]), our approach still shows improvements of 2.1% and 3.7% under the two above feature extractors. While compared to the state-of-the-art DAVE models, our method only underperforms LoCo [6] or CCNet [33] at high tIoU thresholds (0.8 and 0.9), but our overall performance still surpasses that of these models. Combining the experimental results from two tables, our method achieves the highest average mAP under two versions of feature extractors, demonstrating its strong adaptability and robustness.
@@
It is noteworthy that in Table I and Table II, our method performs excellently at low tIoU thresholds (0.5, 0.6, and 0.7), but it is slightly weaker than CCNet [33] and LoCo [6] at tIoU thresholds of 0.8 and 0.9. From the perspective of model structure: (1) CCNet [33] mainly performs temporal modeling at different resolutions, enabling it to focus on information at various temporal granularities; (2) LoCo [6] employs dynamic attention to adjust the temporal scale of the features, which imposes strong constraints on the temporal boundaries of audio-visual events. While our model pay more attention on cross-modal semantic bridging in intermediate layers, enhancing its ability of detection. In summary, CCNet [33] and LoCo [6] are more sensitive to the temporal boundaries of events, with their performance being effective at high tIoU thresholds, whereas our method is more sensitive to the events themselves, thus achieving better performance at low tIoU thresholds.
@@
TABLE III: The ablation study of ESI. $\mathcal{L}_{AVi}$ denotes semantic guided classification loss function in each stage as Eq. (8), ’EF’ denotes the multi-modal early fusion. The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision  at the tIoU thresholds [0.5:0.1:0.9] are reported.
$\mathcal{L}_{AV1}$ $\mathcal{L}_{AV2}$ $\mathcal{L}_{AV3}$ EF 0.5 0.6 0.7 0.8 0.9 Avg.
$\checkmark$ $\checkmark$ $\checkmark$ 58.38 52.90 46.10 37.37 23.24 54.69
$\checkmark$ $\checkmark$ $\checkmark$ 58.50 53.52 46.52 36.94 22.97 54.59
$\checkmark$ $\checkmark$ $\checkmark$ 54.95 49.29 42.98 35.06 24.13 51.82
$\checkmark$ $\checkmark$ 55.69 50.68 44.49 36.44 24.28 52.49
$\checkmark$ 54.72 49.79 43.32 36.20 24.04 51.62
$\checkmark$ $\checkmark$ $\checkmark$ 57.56 51.63 44.48 35.63 21.84 53.50
55.16 49.89 43.80 35.02 21.95 51.61
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 58.61 53.38 46.35 37.12 23.24 54.74
TABLE IV: The ablation study on the weights of the multi-stage semantic guidance. $\alpha_{i}$ denotes the weight of semantic guided classification loss function as in Eq. (8). The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision  at the tIoU thresholds [0.1:0.1:0.9] are reported.
$\alpha_{1},\alpha_{2},\alpha_{3}$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 Avg.
0.5, 0.5, 0.5 72.78 70.47 67.42 63.07 58.65 53.23 45.89 36.69 22.49 54.52
0.3, 0.4, 0.5 72.44 70.30 66.82 62.79 58.11 52.67 46.55 37.39 23.78 54.51
0.3, 0.5, 0.7 72.35 69.98 66.95 63.10 58.32 53.31 46.44 37.45 23.65 54.62
0.3, 0.6, 0.9 72.80 70.59 67.30 63.28 58.61 53.38 46.35 37.12 23.24 54.74
@@
TABLE V: The ablation study on the number of experts in each MoDE layer. Mean Average Precision  at the tIoU thresholds [0.5:0.1:0.9] are reported.
@@
TABLE VI: The ablation study on the layers of MoE in our MoDE module. ’Inf.(sec)’ denotes total inference time in seconds. The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision  at the tIoU thresholds [0.5:0.1:0.9] are reported.
MoE Layers 0.5 0.6 0.7 0.8 0.9 Avg. Inf.(sec)
2 58.48 53.25 46.29 37.07 22.95 54.66 155.67
4 58.61 53.38 46.35 37.12 23.24 54.74 161.04
8 57.78 52.45 46.21 36.70 23.51 54.25 172.75
16 58.66 53.40 46.53 37.49 24.02 54.87 193.32
32 57.89 52.26 45.41 36.73 23.46 54.15 239.54
IV-C Ablation Study

### v2

@@ III-C Early Semantic Interaction
(3) Cross-Modal Pyramid. Our ESI are designed to facilitate progressive cross-modal interaction while gradually reducing modality gaps between two branches, thereby enabling sufficient feature fusion. After the single-modal attention and the audio(visual)-driven mixture, our model achieves audio(visual) temporal aggregation and preliminary cross-modal interaction. To further reduce the modality gap and endow our model with fine-grained temporal boundary awareness (events vary in duration from 0.2s to 70s), we perform directly cross-modal fusion at different temporal resolutions following [1]. Specifically, we temporally downsample the features $\{F_{AV},F_{VA}\}$ with the stride $2^{l_{c}-1}$, where $l_{c}$ is the index of the current pyramid block. There is a cross-attention after a downsampling of audio and visual features. The outputs of the cross-modal pyramid module are $\{\hat{F}_{AV},\hat{F}_{VA}\}\in\mathbb{R}^{T_{l}\times D}$, where $T_{l}$ is the total length of the multi-level pyramid features.
Multi-Stage Semantic Guidance. In this section, we aim to enhance audio-visual event localization by enabling the model to understand event-related content in intermediate layers. However, even with multi-stage feature fusion and temporal modeling, the model still struggles to reduce the semantic gap between audio and visual branches. That is, without explicit semantic constraints in intermediate layers, the model cannot understand and locate event-related segments at early stages. Inspired by [8, 61], we apply an event-related semantic guidance that directly facilitate audio-visual events understanding through multi-stage constraints. Specifically, we employ MLPs on the outputs of Multi-Stage Feature Fusion ($\{\hat{F}_{A},\hat{F}_{V}\}$, $\{F_{AV},F_{VA}\}$, $\{\hat{F}_{AV},\hat{F}_{VA}\}$) for decoding. Each MLP consists of two linear layers with a PReLU activation. Then, we use focal loss functions [62] respectively on the outputs after MLPs for semantic consistency constraints. Note that here we only detect the event categories for each snippet, without regressing the temporal boundaries. The total multi-stage classification loss $\mathcal{L}_{mcls}$ is as:
$\mathcal{L}_{mcls}=\sum_{i=1}^{n}\alpha_{i}(\mathcal{L}_{Ai}+\mathcal{L}_{Vi})$ (8)
where $i\in\{1,2,3\}$, $\alpha_{i}$ is the weight of losses. $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ increase sequentially to ensure the effectiveness of multi-stage optimization from single-modal to multi-modal information. Through these stages of progressive semantic constraints in intermediate layers, Single-Modal Attention can coarsely detect audio and visual events individually, while Audio(Visual)-Driven Mixture and Cross-Modal Pyramid can achieve preliminary and further refined audio-visual event detection, respectively.
@@ III-D Mixture of Dependency Experts
Multi-Modal Temporal Aggregation. Before localizing audio-visual events, we need to aggregate the results of ESI $\{\hat{F}_{AV},\hat{F}_{VA}\}$ and apply temporal modeling to obtain a comprehensive multi-modal representation. By doing this, the model can capture cross-modal long-term relations and filter out irrelevant noise. Specifically, $\hat{F}_{AV}$ and $\hat{F}_{VA}$ are first concatenated together as $Z$. Then we input it into a projection layer to get $Z^{\prime}\in\mathbb{R}^{T_{l}\times C}$ for preliminary fusion while filtering out redundant information, where $C$ is dimension of $Z^{\prime}$, and it is also the number of event categories. Then, to further refine multi-modal information, we employ a temporal modeling for $Z^{\prime}$ as in Eq. (1) and Eq. (2), composed of $N_{1}+N_{2}$ multi-headed self-attention. $\tilde{Z}\in\mathbb{R}^{T_{l}\times C}$is the output after $N_{1}$ temporal attention blocks, and $Z_{t}$ is the final output after $N_{1}+N_{2}$ temporal attention blocks.
Multi-Event Dependencies Extraction. In audio-visual scenarios, there are dependencies between co-occurring events (e.g., thunder is often accompanied by rain and wind). However, these correlations are complex and variable across different scenarios, and the events may have different degrees of overlap. We hope that the model can reveal event dependencies automatically according to different inputs. Mixture-of-Experts model [12, 13, 14], as an approach capable of automatically routing, are well-suited for adaptive complex event dependency processing. Inspired by [63], we further employ a series of MoE layers to handle such issue as shown in the lower right part of Figure 2. Each MoE layer consists of $n$ experts, with only one expert selected at a time. By combining different experts across $m$ layers, complex multi-event dependencies can be revealed adaptively (each MoE layer has $n$ experts, there would be $n^{m}$ possible combinations). Specifically, to ensure that only one expert is selected in each layer at a time, we designed the MoE gate with a Gumbel-Softmax [64]. The output of $i$-th layer is:
@@
where $E_{i,j}$ is the $j$-th expert in the $i$-th layer, $Z_{e}^{i-1}\in\mathbb{R}^{T_{l}\times C}$ is the output of (i1)(i-1)-th MoE layer, $\mathcal{A}$ is initialized as a learnable unit diagonal matrix and adapts to uncover latent dependencies, $C$ is the number of event categories. Finally, the output of the last MoE layer $Z_{e}$ is added with $Z_{t}$ to obtain $\hat{Z}$ for decoding.
@@ III-E Training and Inference
Training. We employ an end-to-end training approach, where three losses are used to constrain the following aspects: a focal loss [62] $\mathcal{L}_{cls}$ for event classification (while addressing class imbalance by focusing on hard examples), a generalized IoU loss [65] $\mathcal{L}_{reg}$ for temporal boundary estimation (while providing scale-invariant regularization and stable gradients for non-overlapping intervals), and $\mathcal{L}_{mcls}$ in Section III-C. The loss functions are expressed as:
@@
$\mathcal{L}_{cls}$ and $\mathcal{L}_{reg}$ are used to constrain the final audio-visual event detection and regression, while $\mathcal{L}_{mcls}$ is used for cross-modal semantic bridging in intermediate layers. More details are provided in the supplementary materials.
Inference. When given all snippets of a video during inference, the model will output candidates that include event categories and temporal boundaries, each accompanied by a confidence score. Then, all candidates are processed by multi-class Soft-NMS [66] to filter out temporal boundaries with high overlap within the same event category.
@@ IV-A Experimental Settings
Datasets. The UnAV-100 dataset proposed by Geng et al. [1] consists of 10,790 videos with a total duration exceeding 126 hours, and includes 30,059 audio-visual events across 100 event categories. In addition to the raw videos, UnAV-100 also provides audio and video features extracted by frozen encoders. It includes two versions: (1) The optical flow features are extracted by RAFT [67], the RGB features are extracted by two-stream I3D [15], while audio features are extracted by VGGish [16]; (2) Audio and visual features are extracted by ONE-PEACE [17], where the visual encoder is fine-tuned on Kinetics-400 [68]. As the standard dataset distribution, the training, validation and testing set are split at a ratio of 3:1:1.
@@ IV-B Comparison with Existing Work
As shown in Table I and Table II, we conduct comprehensive performance comparisons between the proposed method and existing methods. These results show that ESG-Net outperforms other methods in overall performance when using I3D+VGGish [15, 16] or ONE-PEACE [17] for feature extracting. Compared to the existing state-of-the-art TAL models (VSGN [69], TadTR [31], ActionFormer [32] and TridDet [70]), ESG-Net demonstrates significant performance improvements at any threshold. When compared to the baseline method (UnAV [1]), our approach still shows improvements of 2.1% and 4.1% under the two above feature extractors. While compared to the state-of-the-art DAVE models, our method only underperforms LoCo [6] or CCNet [38] at high tIoU thresholds (0.8 and 0.9), but our overall performance still surpasses that of these models. Combining the experimental results from two tables, our method achieves the highest average mAP under two versions of feature extractors, demonstrating its strong adaptability and robustness.
@@
It is noteworthy that in Table I and Table II, our method performs excellently at low tIoU thresholds (0.5, 0.6, and 0.7), but it is slightly weaker than CCNet [38] and LoCo [6] at tIoU thresholds of 0.8 and 0.9. From the perspective of model structure: (1) CCNet [38] mainly performs adaptively temporal modeling at different resolutions, enabling it to focus on information at various temporal granularities; (2) LoCo [6] employs dynamic attention to adjust the temporal scale of the features, which solves the issue of asynchronization in dense audio-visual events detection. In summary, CCNet [38] and LoCo [6] directly enhance the accuracy by imposing strong constraints on the temporal boundaries of audio-visual events. Although our method also achieves significant improvements over the baseline methods, these gains are mainly attributed to the enhanced capability of detecting audio-visual events (thereby mitigating the miss detection). However, since our cross-modal semantic bridging applies audio-visual event constraints instead of audio/visual event constraints in single-modal branches, it inevitably introduces some noise (e.g., the audio-visual event occurs in 3s-5s, the corresponding audio event actually occurs in 2s-6s). Therefore, our method achieves better audio-visual event detection at low tIoU, while exhibiting suboptimal performance at higher tIoU due to the semantic constraints with temporal noise. This can also explain why $\alpha_{i}$ in Eq. (8) needs to increase sequentially (noise in multi-stage semantic constraints is reduced sequentially). In summary, CCNet [38] and LoCo [6] pay more attention on temporal boundaries of audio-visual events, with their performance being effective at high tIoU thresholds, whereas our method is more sensitive to the events themselves, thus achieving better performance at low tIoU thresholds.
TABLE III: Pseudo-label accurancy in AVVP task.
Method Segment-level Event-level Avg.
A V AV Type Event A V AV Type Event
HAN+PPL [71] 62.5 55.3 52.3 56.0 58.3 55.4 51.1 46.9 50.9 50.6 53.9
MA+PPL [71] 61.7 61.8 57.5 60.6 59.4 55.4 57.9 51.6 55.0 52.6 57.4
VALOR [72] 80.4 71.8 63.7 72.0 79.7 72.2 65.9 55.6 64.6 68.0 69.4
UWAV [73] 78.4 74.5 65.5 72.8 78.4 71.1 69.6 57.7 66.1 69.0 70.3
UWAV+ESI 79.2 73.6 65.8 72.8 78.8 72.7 68.9 59.5 67.0 69.9 70.8
@@ A V AV Type Event A V AV Type Event
TABLE IV: The ablation study of ESI. $\mathcal{L}_{AVi}$ denotes semantic guided classification loss function in each stage as Eq. (8), ’ER’ denotes the early multi-modal representations. The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision at the tIoU thresholds [0.5:0.1:0.9] are reported.
$\mathcal{L}_{AV1}$ $\mathcal{L}_{AV2}$ $\mathcal{L}_{AV3}$ ER 0.5 0.6 0.7 0.8 0.9 Avg.
$\checkmark$ $\checkmark$ $\checkmark$ 58.4 52.9 46.1 37.4 23.2 54.7
$\checkmark$ $\checkmark$ $\checkmark$ 58.5 53.5 46.5 36.9 23.0 54.6
$\checkmark$ $\checkmark$ $\checkmark$ 54.9 49.3 43.0 35.1 24.1 51.8
$\checkmark$ 55.0 49.6 43.8 35.3 22.7 51.6
$\checkmark$ 55.7 50.0 43.5 35.6 22.6 52.2
$\checkmark$ 57.0 51.0 44.2 35.7 22.0 53.3
$\checkmark$ 54.7 49.8 43.3 36.2 24.0 51.6
$\checkmark$ $\checkmark$ $\checkmark$ 57.6 51.6 44.5 35.6 21.8 53.5
55.2 49.9 43.8 35.0 22.0 51.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 58.7 53.4 46.7 38.0 24.0 55.1
@@
TABLE V: The ablation study of different structure in ESI. The best mAP in each column is bolded. Mean Average Precision at the tIoU thresholds [0.5:0.1:0.9] are reported.
Method 0.5 0.6 0.7 0.8 0.9 Avg.
Cat 55.1 49.5 43.4 34.6 22.6 51.6
CMPTE(UnAV) 54.8 49.4 42.9 35.1 22.9 51.3
CMCC(CCNet) 54.1 48.8 42.6 34.5 22.6 50.8
Ours 58.7 53.4 46.7 38.0 24.0 55.1
TABLE VI: The ablation study of different structure in MoDE. The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision at the tIoU thresholds [0.5:0.1:0.9] are reported.
Method 0.5 0.6 0.7 0.8 0.9 Avg.
MBM 58.4 53.0 45.7 37.0 22.9 54.5
GCN 58.4 52.6 46.1 36.8 22.4 54.2
Attention 58.0 52.6 45.9 37.1 23.9 54.4
MoS2 Block 58.7 53.2 46.3 37.1 22.0 54.7
TDM(UnAV) 59.2 53.5 46.5 37.4 23.7 55.0
Ours 58.7 53.4 46.7 38.0 24.0 55.1
TABLE VII: The ablation study on the number of experts in each MoDE layer. Mean Average Precision at the tIoU thresholds [0.5:0.1:0.9] are reported.
@@
TABLE VIII: The ablation study on the layers of MoE in our MoDE module. ’Inf.(sec)’ denotes total inference time in seconds. The best mAP in each column is bolded while the second-best is underlined. Mean Average Precision at the tIoU thresholds [0.5:0.1:0.9] are reported.
Layers 0.5 0.6 0.7 0.8 0.9 Avg. Inf.(sec)
2 58.5 53.2 46.3 37.1 23.0 54.7 155.7
4 58.7 53.4 46.7 38.0 24.0 55.1 161.0
8 57.8 52.5 46.2 36.7 23.5 54.3 172.8
16 58.7 53.3 46.5 37.5 24.0 54.9 193.3
32 57.9 52.3 45.4 36.7 23.5 54.2 239.5
IV-C Comparison of Pseudo-Label Accuracy on AVVP
We also extend the multi-modal semantic guidance in intermediate layers of ESI to a similar task, audio-visual video parsing. The experimental results are shown in Table III. UWAV [73] is pre-trained on the UnAV-100 dataset and then evaluated for pseudo-label accuracy on the Look, Listen, and Parse (LLP) dataset [8]. We compute macro F1-scores for three event types: audio (A), visual (V), and audio-visual (AV). Type@AV is the mean of the F1-scores for the A, V, and AV events, while Event@AV is the F1-score of all events in any modality. By integrating multi-stage semantic guidance, "UWAV+ESI" establishes semantic bridging between audio and video modalities within intermediate layers. The results show that "UWAV+ESI" achieves state-of-the-art performance in pseudo-label accuracy on the LLP test set, particularly for audio-visual events, demonstrating both effectiveness and generalizability of ESI.
IV-D Ablation Study